# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

### v2

@@ 3.3 LLM2CLIP: Makes SOTA CLIP more SOTA ever
To compensate for the frozen LLM, and inspired by methods like FuseMix (Vouitsis et al., 2023) and APE (Rosenfeld et al., 2022), we introduce several new linear layers after the LLM as adapters. These layers serve as learnable parameters to improve the alignment between the LLM and CLIP visual encoder. Following the original design of CLIP, we also employ a projector layer to align the dimensions of the two encoders, facilitating the use of CLIP loss for training.
With this powerful LLM-based super text encoder, we achieve a substantial qualitative leap in CLIP’s language comprehension capabilities. The LLM’s open-world knowledge allows the CLIP visual encoder to learn more structured and globally informed visual representations that are aligned with human knowledge. Moreover, this approach enables us to fully utilize high-quality, long, and dense caption datasets without requiring any special architectural adjustments, which previous works like DCI, DreamLip, and Recaption struggled to effectively leverage. As shown in  Table 2, LLM2CLIP makes any existing SOTA CLIP model even more SOTA, significantly surpassing the performance of previous one.
@@ 3.4 Overview and Efficiency Discussion
Here, we highlight some of the design details that significantly improve training efficiency: 1). During the caption contrastive fine-tuning stage, we employ LoRA training for the LLM. Even for a 12B LLM, training with a batch size of 512 only requires around 70GB of GPU memory, making it feasible to run on a single 80GB 8 A100 GPU node. 2). In the LLM2CLIP stage, we freeze the LLM’s gradients and only train the learnable adapter, CLIP’s original Vision encoder, and two projectors. The additional trainable parameters are roughly equivalent to those in the original CLIP, minimizing the overhead. To further reduce the inference cost of using LLMs, we pre-extract all the text features from the training data and store them in memory. This ensures that, even though LLMs provide powerful textual supervision, the memory and computational costs during training remain nearly identical to those of standard CLIP training.
@@ 4.1.1 Caption Contrastive Fine-tuning
We utilized the ShareCaptioner-modified CC-3M dataset (Zheng et al., 2024; Chen et al., 2023), which provides both original captions and augmented dense captions for each image, for contrastive learning. For the Masked Next Token Prediction and caption contrastive fine-tuning stages, we employed the Wikitext-103 dataset (Merity et al., 2016) and the E5 dataset from  Springer et al. (2024) to ensure that the general text domain was not excessively biased.
@@
We trained the language model using LoRA, applying lightweight training with 1 epoch on all datasets. We adopted the average of all LLM output tokens as the text global embedding for a caption. All training parameters follow the design of  BehnamGhader et al. (2024).
@@ 4.1.2 LLM2CLIP Fine-tuning
To compare different configurations, we designed three experimental settings based on dataset size: LLM2CLIP-3M: This configuration only uses the CC-3M dataset for training, representing our lightweight setting. LLM2CLIP-15M: This configuration uses both CC-3M and CC-12M datasets, which is our default setting. All LLM2CLIP models without a specific dataset size label are trained with this 15M data. LLM2CLIP-60M: This configuration scales up the dataset, combining CC-3M, CC-12M, YFCC-15M, and a randomly selected 30M subset from Recaption-1B (Li et al., 2024b). All data used are dense captions rewritten by multimodal large language models (MLLMs). The CC-3M, CC-12M, and YFCC datasets are sourced from  Zheng et al. (2024) using the ShareCaptioner for rewriting, while Recaption data was rewritten using Llama3-LLAVA1.5. We primarily used a mix of original captions and dense captions, with the default mixing ratio being 1:1. LLM2CLIP-S represents a setting where only original data is used for training, matching the original CLIP pretraining distribution to analyze the model’s benefits separately.
@@
All experiments with different datasets were conducted for 4 epochs. We froze the LLM gradients by default. During training, we pre-extracted the text features from the captions using the LLM and stored them in memory to avoid repeated LLM inference, reducing additional computational overhead. We trained the original CLIP vision encoder, the LLM’s learnable adapter, and the projectors for both encoders. Training on the 3M, 15M, and 60M datasets for CLIP ViT-L/14-224 required only 30G per GPU memory and 2, 9, and 45 hours on 8 H100 GPUs, respectively.
4.1.3 Evaluation
Evaluation Dataset.
For short-text datasets, we used the commonly available COCO 2014 5k test set and the Flickr 1k test set as our test datasets. For long-text datasets, we used the 1K subset of the ShareGPT4V (Chen et al., 2023) dataset and the Urban1k (Zhang et al., 2024) dataset, both provided by LongCLIP, along with the DOCCI (Onoe et al., 2024) dataset. The ShareGPT4V-1M dataset was generated using annotations from GPT-4V and ShareCaptioner, with images from Laion, CC, SBU (Ordonez et al., 2011) and MS COCO. We used a randomly selected 1K subset of this dataset. Urban1k consists of captions generated by GPT-4V for 1,000 busy urban view images from Visual Genome. Each caption is a long and complete sentence that describes the image, including types, colors, and relative locations of various attributes. The model can only successfully match the images with the correct captions if it accurately understands and models the detailed attributes in both modalities. The DOCCI dataset contains approximately 1.5K high-resolution images with detailed human-annotated descriptive captions. DOCCI is divided into a training set of 9.6K pairs and a test set of 5.1K pairs. We used the test set for image-lengthy caption retrieval experiments. For the Chinese retrieval tasks, we used the FlickrCN (Lan et al., 2017) and CNCOCO (Li et al., 2018) datasets, which are translated versions of Flickr30K and MS COCO-1K, respectively. These datasets were tested in Chinese. For Llava (Liu et al., 2023) training, we used the standard training and test sets from LLAVA 1.5.
Evaluation Setting.
The experiments in this paper primarily focus on how LLM2CLIP enhances the performance of widely used vanilla CLIP across various dimensions. We mainly compare our approach with the EVA (Fang et al., 2023) and OpenAI CLIP models as baselines, as they are the most widely used SOTA vision encoders in the open-source community. The CLIP models we use for comparison are ViT-B/16222https://huggingface.co/openai/clip-vit-base-patch16, ViT-L/14333https://huggingface.co/openai/clip-vit-large-patch14, and ViT-L/14-336444https://huggingface.co/openai/clip-vit-large-patch14-336. For EVA02 models, we use EVA02 ViT-B/16555https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02_CLIP_B_psz16_s8B.pt, EVA02 ViT-L/14666https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02_CLIP_L_psz14_s4B.pt, and EVA02 ViT-L/14-336777https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02_CLIP_L_336_psz14_s6B.pt, making up a total of six models. For language models, we experimented with four different models: Jina-Embeddings-V2888https://huggingface.co/jinaai/jina-embeddings-v2-base-en, and three popular LLMs from the Llama3 (Dubey et al., 2024) family, namely Llama 3.2 1B, Llama 3 8B, and Mistral-Nemo 12B. To avoid any misunderstanding, the experiments in this paper, unless otherwise noted, use EVA02 ViT-L/14 as the default baseline for comparison. The default language model used is LLaMA 3 8B, trained on a dataset of 15M setting mentioned above.
4.2 Main Results
Directly Replacing CLIP’s Text Encoder with a Vanilla LLM is Harmful.
As shown in the experiments in Table 4, directly replacing the text encoder of EVA02 ViT-L/14 with Llama3-8B significantly degrades retrieval performance. For example, performance on the DOCCI benchmark nearly halves, dropping from 75.0/73.4 to 51.7/50.6. Additionally, the CRA score of the vanilla Llama3-8B is particularly low, indicating that its output features exhibit very poor discriminability for captions. These results clearly show that such an approach imposes a significant burden on CLIP’s learning process.
Table 2: Systematic Comparison Experiment Demonstrating the Performance Improvements of LLM2CLIP.
Methods Flickr30k COCO ShareGPT4V Urban-1k DOCCI
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
ViT-B/16
ALIGN 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1
BLIP 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5
Jina-CLIP 80.6 67.4 55.6 41.1 - - 87.7 88.0 78.7 80.0
Long-CLIP 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4
CLIP 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1
+LLM2CLIP 89.2 78.1 62.2 48.7 98.1 97.4 86.1 90.0 84.1 85.0
EVA02 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0
+LLM2CLIP 88.5 78.0 63.6 49.8 98.0 98.1 84.7 89.7 85.5 86.8
ViT-L/14
Long-CLIP 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6
CLIP 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8
+LLM2CLIP 92.6 81.7 64.9 52.5 98.4 98.4 87.6 92.0 87.6 88.7
EVA02 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0
+LLM2CLIP-3M 89.6 77.3 59.7 48.0 98.3 98.6 87.1 91.1 84.9 87.8
+LLM2CLIP 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4
+LLM2CLIP-30M 92.0 83.5 69.0 55.3 98.9 98.8 93.1 95.0 89.3 91.2
+LLM2CLIP-60M 94.4 83.2 70.4 55.7 99.2 99.4 94.1 95.2 90.2 92.0
ViT-L/14-336
CLIP 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7
+LLM2CLIP 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0
+LLM2CLIP-60M 93.9 82.3 68.5 54.8 98.9 99.1 94.6 95.9 89.6 90.6
EVA02 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4
+LLM2CLIP 93.9 83.8 68.7 55.7 98.8 99.2 89.5 94.2 89.2 91.3
Table 3: Retrieval Performance across Flickr30K-CN and COCO-CN.
@@ 4.2 Main Results
To enhance the discriminability of LLM output features, we applied Caption Contrastive fine-tuning. Llama3-8B-TC and Llama3-8B-CC represent the results of using pure text corpora and a mix of text and augmented CC3M corpora from ShareCaptioner, respectively, both trained with supervised SimCSE loss. As shown in the Table 6, the contrastive learning on mixed caption corpora yields higher CRA scores compared to training on generic text, with noticeable improvements across almost all benchmarks. This difference stems from the fact that LLMs trained on caption-distributed data have stronger discriminability for caption features. These findings highlight the importance of text model discriminability for CLIP training and underscore the rationale behind the caption contrastive fine-tuning in LLM2CLIP.
LLM2CLIP Makes Pretrained SOTA CLIP Even More SOTA.
We applied the LLM2CLIP fine-tuning method to both EVA02 and CLIP models and observed significant performance improvements in Table 2. Even with lightweight fine-tuning, the results substantially surpassed those of the original models pretrained on datasets like Laion2B. Compared to other methods such as LongCLIP and JinaCLIP, which also attempt to fine-tune pretrained CLIP models, our performance gains were transformative. This demonstrates the effectiveness of LLM2CLIP as a method for introducing LLMs to enhance CLIP’s performance.
Table 4: Ablation Study of LLM2CLIP. Here LLM2CLIP-S refers to the results trained on the original short caption dataset.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0
+ Jina-Bert 88.1 77.7 60.5 51.1 83.3 81.0 66.9 68.5 68.9 71.2
       ++ Dense Caption 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9
+ Llama3-8B-S 87.9 75.6 56.7 41.8 55.1 46.1 37.2 35.1 39.3 32.3
       ++ CC Finetuning 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6
        $+$ Dense Caption 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4
LLM’s Built-in Knowledge Directly Benefits CLIP.
In Table 4’s ablation study, we examine the impact of different steps in LLM2CLIP on the EVA02 model. When using only the original captions, there was minimal improvement, as these captions are from the same distribution as those used in EVA02 CLIP’s pretraining. However, we observed that the LLM, through its open-world knowledge, is able to provide a deeper understanding of the original captions, empowering the CLIP model with enhanced capabilities and further boosting performance. Naturally, when dense captions are introduced, the performance improvement becomes even more pronounced.
LLM Improves CLIP’s Ability to Handle Long and Dense Captions.
Thanks to the LLM’s inherent long context windows and strong text understanding capabilities, CLIP can process long caption datasets more effectively without requiring any architectural modifications. This stands in contrast to methods like LongCaption, which require fine-tuning of positional encodings, or Dreamlip, which splits captions into sub-captions. Jina CLIP, for example, enables long-text input by replacing its text encoder with Jina BERT. To verify the LLM’s natural advantage with long texts, we compared our method to Jina CLIP to assess whether the LLM provides a better understanding of long captions. As shown in Table 4, when dense captions are added during training, LLM2CLIP demonstrates significantly greater performance improvements than Jina CLIP, proving that LLMs are better suited to processing long and dense captions.
Larger LLMs Lead to Better Performance.
LLM2CLIP explores the method of replacing CLIP’s text encoder with an LLM and continuing training. In Table X, we thoroughly investigate the performance differences that various versions of LLMs bring to CLIP. In the experiments shown in Table 6, we swapped the text encoder in CLIP with different language models to explore their impact. We maintained the original LLM2CLIP design, with the only change being the initialization of the text encoder. As expected, within the same Llama family, the 8B model significantly outperforms the 1B model. Mistral Namo also shows a slight improvement over the 8B model. It’s important to note that during CLIP training, the LLM’s gradients remain frozen, further proving that larger LLMs, with their richer knowledge, can deliver substantial performance improvements to CLIP. Even so, the 1B LLaMA3.2 model still achieved an impressive performance boost for EVA02 CLIP, with an average improvement of 11.1% — an impressive feat considering that EVA02 is already one of the SOTA open-source CLIP models.
Ratio Flickr30k COCO ShareGPT4v Urban-1k DOCCI
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
100% 85.5 72.7 60.1 46.9 98.7 99.0 88.7 93.9 88.0 90.5
75% 92.4 82.6 68.5 54.2 98.7 99.3 89.0 94.3 88.1 90.2
50% 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4
25% 93.0 82.8 68.1 54.8 98.4 98.7 87.7 92.9 87.9 90.0
0% 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6
Table 6: Comparison of various text encoders.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average CRA
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
EVA02 Vit-L/14 89.8 73.3 63.8 63.8 89.3 91.9 68.5 73.3 75.0 73.4 76.2 69.8
+Jina Bert 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 78.2 74.2
+Llama3-8B 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 66.3 18.4
+Llama3-8B-TC 92.7 82.1 68.1 54.6 97.7 98.2 88.9 93.8 85.0 87.8 84.8 71.3
+Llama3-8B-CC 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 85.6 73.0
+Llama3.2-1B-CC 91.6 81.3 65.8 52.5 98.3 98.2 84.5 91.9 83.4 86.4 83.4 72.8
+Mistral-Nemo-12B-CC 93.5 83.7 68.5 54.7 98.6 98.9 90.4 94.3 88.0 89.7 86.0 73.3
Table 7: Performance of Llava 1.5. The best results are highlighted in bold.We explored whether LLM2CLIP could enhance complex image understanding tasks by modifying Llava’s visual encoder.

### v3

@@ 3.2 LLM2CLIP: Use LLM to Finetuning
To compensate for the frozen LLM, and inspired by methods like FuseMix [41] and APE [35], we introduce several new linear layers after the LLM as adapters. These layers serve as learnable parameters to improve the alignment between the LLM and CLIP visual encoder. Following the original design of CLIP, we also employ a projector layer to align the dimensions of the two encoders, facilitating the use of CLIP loss for training.
With this powerful LLM-based super text encoder, we achieve a substantial qualitative leap in CLIP’s language comprehension capabilities as results in Table 2. The LLM’s open-world knowledge allows the CLIP visual encoder to learn more structured and globally informed visual representations that are aligned with human knowledge. Moreover, this approach enables us to fully utilize high-quality, long, and dense caption datasets without requiring any special architectural adjustments, which previous works like DCI, DreamLip, and Recaption struggled to effectively leverage.
@@ 3.3 Overview and Efficiency Discussion
Here, we highlight some of the design details that significantly improve training efficiency: 1). During the caption contrastive fine-tuning stage, we employ LoRA training for the LLM. Even for a 12B LLM, training with a batch size of 512 only requires around 70GB of GPU memory, making it feasible to run on a single 80GB 8 A100 GPU node. 2). In the LLM2CLIP stage, we freeze the LLM’s gradients and only train the learnable adapter, CLIP’s original Vision encoder, and two projectors. The additional trainable parameters are roughly equivalent to those in the original CLIP, minimizing the overhead. To further reduce the inference cost of using LLMs, we pre-extract all the text features from the training data and store them in disk memory. This ensures that, even though LLMs provide powerful textual supervision, the memory and computational costs during training remain nearly identical to those of standard CLIP training.
@@
Methods Flickr30k COCO ShareGPT4V Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
ViT-B/16
ALIGN 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
BLIP 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
Jina-CLIP 80.6 67.4 55.6 41.1 - - 87.7 88.0 78.7 80.0 75.7 69.1
Long-CLIP 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
CLIP 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
+LLM2CLIP 89.2 78.1 62.2 48.7 98.1 97.4 86.1 90.0 84.1 85.0 83.9 79.8
EVA02 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0 74.0 65.6
+LLM2CLIP 88.5 78.0 63.6 49.8 98.0 98.1 84.7 89.7 85.5 86.8 84.1 80.5
ViT-L/14
Long-CLIP 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
CLIP 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
+LLM2CLIP 92.6 81.7 64.9 52.5 98.4 98.4 87.6 92.0 87.6 88.7 86.2 82.7
EVA02 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
+LLM2CLIP-3M 89.6 77.3 59.7 48.0 98.3 98.6 87.1 91.1 84.9 87.8 83.9 80.6
+LLM2CLIP 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
+LLM2CLIP-30M 92.0 83.5 69.0 55.3 98.9 98.8 93.1 95.0 89.3 91.2 88.5 84.8
+LLM2CLIP-60M 94.4 83.2 70.4 55.7 99.2 99.4 94.1 95.2 90.2 92.0 89.7 85.1
ViT-L/14-336
CLIP 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7 74.4 62.2
+LLM2CLIP 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0 86.6 83.3
+LLM2CLIP-60M 93.9 82.3 68.5 54.8 98.9 99.1 94.6 95.9 89.6 90.6 89.1 84.5
EVA02 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4 79.3 72.3
+LLM2CLIP 93.9 83.8 68.7 55.7 98.8 99.2 89.5 94.2 89.2 91.3 88.0 84.8
Table 2: Systematic Comparison Experiment Demonstrating the Performance Improvements of LLM2CLIP. Results with a blue background represent the outcomes of LLM2CLIP.
@@ 4.1.1 Caption Contrastive Fine-tuning
For contrastive learning, we utilized the ShareCaptioner-modified CC-3M dataset [45, 5], which includes both original captions and augmented dense captions for each image. In addition, we used the Wikitext-103 dataset [28] and the E5 dataset [37] during the Masked Next Token Prediction and caption contrastive fine-tuning stages to ensure coverage of the general text domain without excessive bias.
@@
The language model was trained with LoRA for lightweight training, completing one epoch across all datasets. We used the average of all LLM output tokens as the global text embedding for a caption. All training parameters followed the configuration detailed in [3]. The design of the Adaptor is the same as FuseMix [41], using a simple inverted bottleneck MLP architecture.
@@ 4.1.2 LLM2CLIP Fine-tuning
We defined three experimental configurations based on dataset size:
  • LLM2CLIP-3M: This configuration uses only the CC-3M dataset, serving as our lightweight setting.
  • LLM2CLIP-15M: Our default configuration, combining CC-3M and CC-12M datasets.
  • LLM2CLIP-60M: A larger configuration that combines CC-3M, CC-12M, YFCC-15M, and a 30M randomly selected subset from Recaption-1B [20].
All datasets included dense captions rewritten by multimodal large language models (MLLMs). Specifically, the CC-3M, CC-12M, and YFCC datasets were sourced from [45] and rewritten using ShareCaptioner, while Recaption data were rewritten with Llama3-LLAVA1.5. A 1:1 mix of original captions and dense captions was used by default.
@@
Each configuration was trained for four epochs. We froze the LLM gradients to reduce computational overhead and pre-extracted text features using the LLM, storing them on disk to avoid repeated inference during training. Inspired by Fusemix’s efficient training strategy, we added a four-layer linear adapter after the LLM and trained with gradients enabled for the CLIP ViT model to better integrate LLM knowledge. Additionally, we introduced two projector layers after each encoder for dimensional alignment.
Training the CLIP ViT-L/14-224 model required approximately 30GB of GPU memory per GPU and took 2, 9, and 45 hours on 8 H100 GPUs for the 3M, 15M, and 60M datasets, respectively. We employed the AdamW [26] optimizer, utilizing a total batch size of 4096, and set the learning rate to 1e-5 with a cosine decay rate of 0.05.
For language models, we tested four configurations: Jina-Embeddings-V2, and three popular models from the Llama3 family [8]: Llama 3.2 1B, Llama 3 8B, and Mistral-Nemo 12B. Unless otherwise noted, experiments used EVA02 ViT-L/14 as the baseline and Llama 3 8B as the default language model.
@@
Table 3: Retrieval Performance across Flickr30K-CN and COCO-CN.
Methods Imagenet
0-shot Linear
CLIP L/14-336 76.6 84.8
+ LLM2CLIP 74.1 85.2
CLIP L/14 75.5 83.9
+ LLM2CLIP 71.8 84.4
CLIP B16 70.2 79.9
+ LLM2CLIP 67.3 80.6
Table 4: Imagenet 0-shot classification and linear probe results.
@@
Table 5: Performance of Llava 1.5 benchmarks. The best results are highlighted in bold.We explored whether LLM2CLIP could enhance complex image understanding tasks by modifying Llava’s visual encoder.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average CRA
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5 27.1
+Jina Bert 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 79.5 77.0 29.68
+Llama3-8B 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 68.6 64.0 5.22
+Llama3-8B-TC(LLM2Vec) 92.7 82.1 68.1 54.6 97.7 98.2 88.9 93.8 85.0 87.8 86.5 83.3 28.42
+Llama3-8B-CC 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2 29.46
+Llama3.2-1B-CC 91.6 81.3 65.8 52.5 98.3 98.2 84.5 91.9 83.4 86.4 84.7 82.1 29.42
+Mistral-Nemo-12B-CC 93.5 83.7 68.5 54.7 98.6 98.9 90.4 94.3 88.0 89.7 87.8 84.3 29.96
Table 6: Comparison of various text encoders.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
+ Jina-Bert 88.1 77.7 60.5 51.1 83.3 81.0 66.9 68.5 68.9 71.2 73.5 70.0
       ++ Dense Caption 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 79.5 77.0
+ Llama3-8B 87.9 75.6 56.7 41.8 55.1 46.1 37.2 35.1 39.3 32.3 55.2 46.2
       ++ Dense Caption 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 68.6 64.0
+ Llama3-8B-CC 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6 83.4 80.3
       ++ Dense Caption 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
Table 7: Ablation analysis of performance variations between short simple captions and long & dense captions using different text encoders. We conducted experiments on the EVA02 ViT-L/14 model, replacing its original text encoder with Jina-BERT (+ Jina-BERT), the original Llama3 (+ Llama3-8B), and Llama3 fine-tuned with Caption Contrastive learning (+ Llama3-8B-CC). Fine-tuning was performed on datasets containing only short simple captions and datasets with a 50% probability mix of long dense captions. The objective of this experiment was to isolate the effects of LLMs and long dense captions, and to demonstrate that the design of LLM2CLIP enhances the efficiency of dense caption utilization. The results with a red background indicate that the LLM without fine-tuning has a significant negative effect on CLIP.
Ratio Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
100% 85.5 72.7 60.1 46.9 98.7 99.0 88.7 93.9 88.0 90.5 84.2 80.6
75% 92.4 82.6 68.5 54.2 98.7 99.3 89.0 94.3 88.1 90.2 87.3 84.1
50% 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
25% 93.0 82.8 68.1 54.8 98.4 98.7 87.7 92.9 87.9 90.0 87.0 83.8
0% 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6 83.4 80.3
Adaptor Layers Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
0 91.3 79.9 64.0 50.2 97.1 97.7 88.1 90.2 81.0 86.1 84.3 80.8
1 92.1 80.6 65.9 51.8 97.8 98.0 90.9 91.8 84.3 87.9 86.2 82.0
4 92.1 79.9 66.0 52.2 97.4 98.0 93.0 92.1 85.0 88.0 86.7 82.0
Table 9: Performance comparison across benchmarks (Flickr30k, COCO, ShareGPT4v, Urban-1k, and DOCCI) using different adaptor layer numbers. Results are reported for image-to-text (I2T) and text-to-image (T2I) retrieval tasks. The best result in each column is highlighted in bold. All experiments were conducted using OpenAI’s CLIP model with a ViT architecture at 224 resolution, trained on 15M data points.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
Bidirectional Attention + MNTP + Sup. SimCLR 83.3 73.5 55.3 45.6 98.1 98.0 81.5 87.9 78.8 82.6 79.4 77.5
Bidirectional Attention + Sup. SimCLR 84.6 74.5 55.6 46.2 98.0 97.6 79.3 88.5 80.0 82.6 79.5 77.9
Sup. SimCLR 82.0 73.5 54.6 44.8 97.6 97.6 82.2 87.9 76.9 81.5 78.7 77.1
Table 10: Performance comparison across benchmarks (Flickr30k, COCO, ShareGPT4v, Urban-1k, and DOCCI) using different methods. Results are reported for image-to-text (I2T) and text-to-image (T2I) retrieval tasks. The best result in each column is highlighted in bold. The experiments were conducted using the Llama 3.2 1B and the EVA02 ViT-L/14 model on 3M data. It can be observed that MNTP has no significant effect, while bidirectional attention provides a slight performance improvement.
4.2 Main Results
Evaluation Setting.
Our experiments assessed LLM2CLIP’s ability to enhance vanilla CLIP performance across multiple dimensions. We compared our approach against state-of-the-art (SOTA) models, including EVA [10] and OpenAI CLIP. Specifically, we evaluated CLIP ViT-B/16, ViT-L/14, and ViT-L/14-336, along with EVA02 ViT-B/16, EVA02 ViT-L/14, and EVA02 ViT-L/14-336, yielding a total of six comparison models.
4.2.1 Short- and Long-Text Retrieval
Evaluation Dataset.
For short-text retrieval, we used the MSCOCO [23] 5K test set and the Flickr [43] 1K test set. For long-text retrieval, we employed datasets from LongCLIP, including a 1K subset of ShareGPT4V [5], the Urban1K [44] dataset, and the DOCCI [29] dataset.
The ShareGPT4V-1M dataset consists of captions generated using GPT-4V and ShareCaptioner, covering images from LAION, CC, SBU [30], and MS COCO. Urban1K includes captions for 1,000 urban scene images, each richly annotated with detailed descriptions. DOCCI contains 1.5K high-resolution images with human-annotated captions and was used for retrieval evaluation.
LLM2CLIP Improves CLIP for Long- and Short-Text Retrieval.
As shown in Table 2, LLM2CLIP fine-tuning yielded significant performance improvements, even surpassing models pretrained on large datasets like LAION2B. Compared to LongCLIP and JinaCLIP, our method achieved transformative gains, showcasing its efficiency in leveraging LLM knowledge to enhance pretrained CLIP models.
4.2.2 Cross-Lingual Retrieval
Evaluation Dataset.
For Chinese retrieval tasks, we used FlickrCN [18] and CNCOCO [21], which are translated versions of Flickr30K and MS COCO-1K, respectively. LLAVA training and evaluation were conducted using the standard splits from LLAVA 1.5 [24].